OpenStudy (anonymous):

how do factor a^3+125 using the difference of two cubes

OpenStudy (anonymous):

it is not the difference of two cubes,it is the sum of two cubes \[a^2+b^2=(a+b)(a^2-ab+b^2)\] put \[a=a, b=5\]

OpenStudy (anonymous):

Technically, it can be a difference: \[a^3+125=a^3-(-5)^3\] but that's just semantics.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

this is what the answer is (a+5)(a^2-5a+25)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

I need to show the steps of getting that answer